# Tabular foundation model to detect empathy from visual cues

### v1

@@ 3.3. In-context learning with tabular foundation model
(4) $p(y=1|\mathbf{x},\mathcal{D}_{\text{train}};\theta)=f_{\theta}(\mathbf{x},\mathcal{D}_{\text{train}})$
Tabular foundation models such as TabPFN and TabICL support in-context learning. These foundation models are pre-trained on millions of synthetic tabular datasets to predict labels for test samples given a set of labelled training samples as context. Formally, they approximate the posterior predictive distribution:
(5) $p(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})\approx q_{\theta}(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$
@@
(6) $\hat{y}=\arg\max_{y\in\{0,1\}}q_{\theta}(y|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$
A simplified representation of the TabPFN architecture is presented in Figure 2 (refer to (Hollmann et al., 2023, 2025) for details). The model includes two separate encoders (X and Y) for processing features and labels, respectively. These encoders also perform basic preprocessing such as data cleaning and input normalisation. The encoded features and labels are concatenated and passed through 12 transformer blocks. Each block includes two multi-head attention mechanisms: one for feature-wise attention and another for item-wise attention. Finally, a decoder transforms the output using a linear layer followed by GELU activation.
3.4. Fine-tuning tabular foundation model
While tabular foundation models provide strong in-context learning performance, their performance can be further enhanced by fine-tuning using the training data. We update the model parameters $\theta$ to $\theta^{\prime}$ by minimising the following binary cross entropy loss with sigmoid activation and temperature scaling:
(7) $\mathcal{L}_{\text{fine}}=-\frac{1}{N_{\text{fine}}}\sum_{i=1}^{N_{\text{fine}}}\left[y_{i}\cdot\log(\sigma(z_{i}^{+}/\tau))+(1-y_{i})\cdot\log(1-\sigma(z_{i}^{+}/\tau))\right]$
@@ 3.4. Fine-tuning tabular foundation model
To prevent overfitting during fine-tuning, we employ adaptive early stopping and also save the best model based on validation performance.
3.5. Evaluation protocol
@@
(8) $\begin{split}\exists(\mathbf{x}_{i},y_{i}),&(\mathbf{x}_{j},y_{j})\in\mathcal{D}:\\ \text{subject}(\mathbf{x}_{i})=&\text{subject}(\mathbf{x}_{j})\land(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(f)}\land(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(f)}\end{split}$
@@
(9) $\mathcal{D}_{\text{train}}^{(j)}=\mathcal{D}\setminus\mathcal{D}_{s_{j}},\quad\mathcal{D}_{\text{test}}^{(j)}=\mathcal{D}_{s_{j}}$
@@
(10) $\forall(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(j)},(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(j)}:\text{subject}(\mathbf{x}_{i})\neq\text{subject}(\mathbf{x}_{j})$
@@
4. Experiments
4.1. Dataset
Mathur et al. (2021) released the human-robot interaction (HRI) dataset, which consists of 122 videos (total of 6.9 hours) of interactions between 42 subjects and a robot storyteller. Each interaction followed a controlled protocol where participants listened to three distinct stories, each lasting approximately three minutes. The stories were delivered with randomly assigned narrative voice conditions: half in a first-person narrative voice where the robot told stories about itself, and half in a third-person narrative voice where the robot narrated from an external perspective. After each story, participants completed an 8-item questionnaire using a Likert scale, which produced the ground truth empathy score for each interaction. The labels were binarised – “empathic” and “less-empathic” – using the median empathy score as the threshold. The label distribution is balanced.
For each video, 709 visual behavioural features were extracted using the OpenFace 2.2.0 toolkit (Amos et al., 2016), which captured comprehensive facial and gaze information from human subjects as they listened to the robot’s stories. These features include eye gaze directions, intensity and presence of 17 facial action units, facial landmarks, head pose coordinates and point-distribution model parameters representing facial location, scale, rotation and deformation.
Table 1. Hyperparameter tuning results of 50 trials in Optuna, optimised for 5-fold 10-repeats stratified cross-validated average classification accuracy. Hyperparameters corresponding to the best accuracy are highlighted in bold.
@@ 4.1. Dataset
Table 2. 5-fold 10-repeat stratified cross-validation (average $\pm$ standard deviation) scores of tabular foundation models versus baseline models. Single scores refer to the scores reported in the literature. Highest scores are formatted in bold.
@@
XGBoost(Mathur et al., 2021)a 0.69 0.72 0.69 0.69 0.69
XGBoost(Mathur et al., 2021)b $0.598\pm 0.071$ $0.623\pm 0.099$ $0.604\pm 0.079$ $0.589\pm 0.135$ $0.588\pm 0.090$
@@
LSTM (Mathur et al., 2021) 0.65 0.71 0.65 0.70
TCN (Mathur et al., 2021) 0.54 0.54 0.76 0.62
TabICL in-context learning $0.634\pm 0.084$ $0.665\pm 0.103$ $0.637\pm 0.100$ $0.660\pm 0.139$ $0.639\pm 0.094$
TabPFN in-context learning $0.622\pm 0.091$ $0.653\pm 0.100$ $0.621\pm 0.096$ $0.651\pm 0.148$ $0.627\pm 0.103$
TabPFN fine-tuned $\textbf{0.715}\pm 0.073$ $0.700\pm 0.091$ $\textbf{0.773}\pm 0.126$ $0.659\pm 0.156$ $\textbf{0.691}\pm 0.096$
@@
    Scores reported in (Mathur et al., 2021) without specification of metric type (average or best).
@@
    Our re-implementation based on reported hyperparameters; no code was released for (Mathur et al., 2021).
@@
Table 3. Generalisation on unseen subjects through leave-one-subject-out cross-validation. Highest scores are highlighted in bold.
@@
TabICL in-context learning 0.541 0.519 0.547 0.475 0.509
TabPFN in-context learning 0.459 0.472 0.451 0.377 0.411
TabPFN fine-tuned (all layers trainable) 0.689 0.642 0.780 0.525 0.627
TabPFN fine-tuned (Frozen X and Y encoders) 0.730 0.669 0.818 0.590 0.686
4.2. Implementation details
We use PyTorch 2.2.0, scikit-learn 1.6.1 and other libraries in Python 3.10.12 to implement our codebase. Details of other libraries and their versions are provided in our GitHub repository. Experiments are run on a single AMD Instinct™ MI250X GPU.
Hyperparameters are tuned using Optuna (Akiba et al., 2019) with its default sampler (tree-structured Parzen estimator) and pruner (median stopping rule). For a fair comparison, we optimise two key hyperparameters for 50 trials for all baseline models and TabPFN fine-tuning setup, with an objective of maximising the average classification accuracy across 5-fold 10-repeats stratified cross-validation. We did not tune the hyperparameters of TabICL and TabPFN in the ICL setup; instead, we used their default configurations. Table 1 presents the list of hyperparameters and corresponding models. The same set of selected hyperparameters is then applied in the leave-one-subject-out experiments to experiment generalisation to unseen subjects.
Tabular foundation models. We use the official pre-trained model release of TabPFN v2222https://huggingface.co/Prior-Labs/TabPFN-v2-clf and TabICL [^3] https://huggingface.co/jingang/TabICL-clf foundation models. Since TabICL’s training code remains unreleased as of 07 April 2025, we use it in the in-context learning experiment only, whereas we experiment with TabPFN in both in-context learning and fine-tuning settings.
Baseline models. We benchmark against XGBoost, which Mathur et al. (2021) reported as the best model for the HRI dataset. Since no official implementation of their XGBoost usage was released, we re-implemented their XGBoost configuration using the three hyperparameters specified in (Mathur et al., 2021). We use default values of other hyperparameters unspecified in (Mathur et al., 2021). We denote this version as XGBoost(Mathur et al., 2021) throughout the paper. Since our implementation of XGBoost(Mathur et al., 2021) failed to reproduce the performance metrics reported in (Mathur et al., 2021), presumably due to undisclosed hyperparameter or training details, we implemented an additional XGBoost model with hyperparameter optimisation via Optuna. Apart from XGBoost, we further implemented random forest (RF) and support vector machine (SVM) algorithms with their default hyperparameters as baseline models.
4.3. Results: Competitive performance on established evaluation protocol
Following the 5-fold stratified 10-repeats cross-validation evaluation protocol established by Mathur et al. (2021), Table 2 compares cross-validated average and standard deviation scores of in-context learning and fine-tuning of tabular foundation models against the baseline models as well as the performance reported in the literature. Since TabPFN’s recommended input capacity is up to 500 features (Hollmann et al., 2025), we select 500 features. For a fair comparison, we also select 500 features for baseline models, including XGBoost, RF and SVM.
Tabular foundation models in ICL setup surpass or achieve comparable performance to baseline models across all evaluation metrics. Assuming the scores reported for XGBoost(Mathur et al., 2021) are average across 50 folds, the fine-tuned TabPFN still outperforms it in classification accuracy, precision and F1 scores. The discrepancy between the reported scores for XGBoost(Mathur et al., 2021) and our re-implementation is likely due to undisclosed hyperparameters or training details.
TabICL outperforms TabPFN. Consistent with the findings of (Qu et al., 2025), our experiment confirms the superior performance of TabICL over TabPFN in the ICL setting. Note that we did not perform any hyperparameter tuning for the ICL experiments.
Fine-tuning tabular foundation model further improves performance. When fine-tuned, TabPFN shows a substantial performance gain and surpasses TabICL. It is worthwhile to note that this improvement comes at the cost of additional computation.
Tabular foundation models outperform other deep learning approaches. As shown in Table 2, tabular foundation models outperform LSTM and TCN deep learning architectures. Their lower performance may be attributed to the use of raw temporal sequences as input, rather than statistical attributes.
@@ 4.3. Results: Competitive performance on established evaluation protocol
Table 4. TabPFN fine-tuning results (5-fold 10-repeats stratified subject-unaware cross-validation) with different experimental settings. Bold indicates the selected configuration.
@@
4.4. Results: Improved generalisation on unseen subjects
Given the presence of multiple samples per subject (122 samples from 45 subjects), we adopt a train-test split that accounts for subject identity. Specifically, we use leave-one-subject-out cross-validation, where the splitting is based on participant ID available on the dataset. As this approach introduces no randomness, the results reported in Table 3 are deterministic and presented without any standard deviation.
Fine-tuning gives a significant boost to TabPFN cross-subject generalisation. Its classification accuracy improved from 0.459 to 0.730 after fine-tuning. Although the model had no access to the data from the test subject, it effectively learned patterns from other subjects, which suggests good cross-subject generalisation.
4.5. Ablation study
@@
We report a series of experiments on TabPFN fine-tuning (Table 4). Key hyperparameters explored include learning rate, batch size and optimiser choice. We tried the recent scheduler-free optimiser (Defazio et al., 2024) as well as the standard AdamW optimiser, with and without a learning rate scheduler. Since catastrophic forgetting is a challenge when fine-tuning pre-trained models, we tried a warm-up strategy (where the learning rate is linearly increased from a small initial value) with the AdamW optimiser. Specifically, we used a one-cycle learning rate scheduler with cosine annealing. Overall, TabPFN showed minimal performance variations across these optimiser choices, with AdamW and one-cycle scheduler showing the best results. Like optimiser choice, batch size has a modest effect on performance. Increasing it from 8 to 32 improved classification accuracy from 0.701 to 0.708 and AUC from 0.689 to 0.696.
We also tested freezing some pre-trained layers of TabPFN. Freezing the X and Y encoders led to improved performance and faster fine-tuning, so we retained them as frozen and keeping the remaining layers trainable.
@@
We further varied the softmax temperature and the number of input features. Using 500 features led to a slight performance improvement. A softmax temperature of 1.0 (i.e., no temperature scaling) yielded the best classification accuracy (Table 4). Based on these findings, the final model uses the AdamW optimiser with a one-cycle scheduler, a batch size of 32, frozen X and Y encoders, a softmax temperature of 1.0, and 500 selected input features.
@@
Figure 3. In-context learning performance of TabICL and TabPFN across different proportions of in-context data (i.e., training data used in the single forward pass of in-context learning). Results are based on 5-fold, 10-repeat stratified cross-validation without subject group separation.
@@
Figure 3 shows the in-context learning performance of TabICL and TabPFN across varying amounts of context data. In the ICL setup, the context data – i.e., training data – is fed into the model during the single forward pass. Both models generally exhibit improved performance as the amount of context data increases, which suggests the potential for further gains with additional data. However, TabPFN’s AUC peaked when using 80% of the available context data.
4.6. Limitations

### v2

@@ III-D In-context learning with tabular foundation model
$p(y=1|\mathbf{x},\mathcal{D}_{\text{train}};\theta)=f_{\theta}(\mathbf{x},\mathcal{D}_{\text{train}})$ (4)
TFMs such as TabPFN and TabICL, by design, support ICL. These foundation models are pre-trained on millions of synthetic tabular datasets to predict labels for test samples, given a set of labelled training samples as context. Formally, they approximate the posterior predictive distribution:
$p(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})\approx q_{\theta}(y_{\text{test}}|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$ (5)
@@
$\hat{y}=\arg\max_{y\in\{0,1\}}q_{\theta}(y|\mathbf{x}_{\text{test}},\mathcal{D}_{\text{train}})$ (6)
A simplified representation of the TabPFN architecture is included in Figure 2 (refer to [12, 13] for details). The model includes two separate encoders (X and Y) for processing features and labels, respectively. These encoders also perform basic preprocessing such as data cleaning and input normalisation. The encoded features and labels are concatenated and passed through 12 transformer blocks. Each block includes two multi-head attention mechanisms: one for feature-wise attention and another for item-wise attention. Finally, a decoder transforms the output using a linear layer followed by GELU activation.
III-E Fine-tuning tabular foundation model
While TFMs provide strong ICL performance, their performance can be further enhanced by fine-tuning using the training data. We update the model parameters $\theta$ to $\theta^{\prime}$ by minimising the following binary cross-entropy loss with sigmoid activation and temperature scaling:
$\mathcal{L}_{\text{fine}}=-\frac{1}{N_{\text{fine}}}\sum_{i=1}^{N_{\text{fine}}}\Bigl{[}y_{i}\cdot\log(\sigma(z_{i}^{+}/\tau))\\ +(1-y_{i})\cdot\log(1-\sigma(z_{i}^{+}/\tau))\Bigr{]}$ (7)
@@ III-E Fine-tuning tabular foundation model
Catastrophic forgetting is a common challenge with fine-tuning pre-trained models. To address this, we apply a warm-up strategy (where the learning rate is linearly increased from a small initial value) with the optimiser. Specifically, we used a one-cycle learning rate scheduler with cosine annealing. To prevent overfitting during fine-tuning, we employ adaptive early stopping, which dynamically adjusts the patience parameter (i.e., the number of steps before stopping the training) throughout training rather than relying on a fixed value. We save the best model based on validation performance.
III-F Evaluation protocol
@@
$\begin{split}\exists(\mathbf{x}_{i},y_{i}),&(\mathbf{x}_{j},y_{j})\in\mathcal{D}:\\ \text{subject}(\mathbf{x}_{i})=&\text{subject}(\mathbf{x}_{j})\land(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(f)}\land(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(f)}\end{split}$ (8)
@@
$\mathcal{D}_{\text{train}}^{(j)}=\mathcal{D}\setminus\mathcal{D}_{s_{j}},\quad\mathcal{D}_{\text{test}}^{(j)}=\mathcal{D}_{s_{j}}$ (9)
@@
$\forall(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{\text{train}}^{(j)},(\mathbf{x}_{j},y_{j})\in\mathcal{D}_{\text{test}}^{(j)}:\text{subject}(\mathbf{x}_{i})\neq\text{subject}(\mathbf{x}_{j})$ (10)
@@
IV Experiments and Results
IV-A Dataset
[1] released the human-robot interaction (HRI) dataset, which consists of 122 videos (a total of 6.9 hours) of interactions between 42 subjects and a robot storyteller. Each interaction followed a controlled protocol where participants listened to three distinct stories, each lasting approximately three minutes. The stories were delivered with randomly assigned narrative voice conditions: half in a first-person narrative voice where the robot told stories about itself, and half in a third-person narrative voice where the robot narrated from an external perspective. After each story, participants completed an 8-item questionnaire using a Likert scale, which produced the ground truth empathy score for each interaction. [1] binarised the labels into “empathic” and “less-empathic” categories using the median empathy score as a threshold. The label distribution is balanced.
For each video, 709 visual behavioural features were extracted using the OpenFace 2.2.0 toolkit [35], which captured comprehensive facial and gaze information from human subjects as they listened to the robot’s stories. These features include eye gaze directions, intensity and presence of 17 facial action units, facial landmarks, head pose coordinates and point-distribution model parameters representing facial location, scale, rotation and deformation.
TABLE I: Hyperparameter tuning results of 50 trials in Optuna, optimised for 5-fold 10-repeats stratified cross-validated average classification accuracy. Hyperparameters corresponding to the best accuracy are highlighted in bold.
@@ IV-A Dataset
TABLE II: 5-fold 10-repeat stratified cross-validation (average $\pm$ standard deviation) scores of our proposed TFMPathy versus baseline models. Single scores refer to the scores reported in the literature. Highest scores are formatted in bold. ICL refers to in-context learning, and FT refers to fine-tuning.
@@
XGBoost[1]a 0.69 0.72 0.69 0.69 0.69
XGBoost[1]b $0.598\pm 0.071$ $0.623\pm 0.099$ $0.604\pm 0.079$ $0.589\pm 0.135$ $0.588\pm 0.090$
@@
LSTM [1] 0.65 0.71 0.65 0.70
TCN [1] 0.54 0.54 0.76 0.62
TFMPathyICL(TabICL) $0.634\pm 0.084$ $0.665\pm 0.103$ $37\pm 0.100 0.660$$60\pm 0.139$ $0.639\pm 0.094$
TFMPathyICL(TabPFN) $0.622\pm 0.091$ $0.653\pm 0.100$ $0.621\pm 0.096$ $0.651\pm 0.148$ $0.627\pm 0.103$
TFMPathyFT(TabPFN) $\textbf{0.715}\pm 0.073$ $0.700\pm 0.091$ $\textbf{0.773}\pm 0.126$ $0.659\pm 0.156$ $\textbf{0.691}\pm 0.096$
@@
    Scores reported in [1] without specification of metric type (average or best).
@@
    Our re-implementation based on reported hyperparameters; no code was released for [1].
@@
IV-B Implementation details
We use PyTorch 2.2.0, scikit-learn 1.6.1 and other libraries in Python 3.10.12 for experiments. Details of other libraries and their versions are provided in our public GitHub repository. Experiments are run on a single AMD Instinct™ MI250X GPU (64 GB GCD).
Hyperparameters are tuned using Optuna [36] with its default sampler (tree-structured Parzen estimator) and pruner (median stopping rule). For a fair comparison, we optimise two key hyperparameters for 50 trials for all baseline models and TabPFN fine-tuning setup, with an objective of maximising the average classification accuracy across 5-fold 10-repeats stratified cross-validation. We did not tune the hyperparameters of TabICL and TabPFN in the ICL setup; instead, we used their default configurations. Table I presents the list of hyperparameters and corresponding models. The same set of selected hyperparameters is then applied in the leave-one-subject-out experiments to experiment with generalisation to unseen subjects.
IV-B1 Tabular foundation models
We use the official pre-trained release of TabPFN v2333https://huggingface.co/Prior-Labs/TabPFN-v2-clf and TabICL [^4] https://huggingface.co/jingang/TabICL-clf foundation models. Since TabICL’s training code was unavailable, we use it in the ICL experiment only, whereas we experiment with TabPFN in both ICL and fine-tuning settings.
IV-B2 Baseline models
We benchmark against XGBoost, which [1] reported as the best model for the HRI dataset. Since no official implementation of their XGBoost usage was released, we re-implemented their XGBoost configuration using the three hyperparameters specified in [1]. We use the default values of other hyperparameters unspecified in [1]. We denote this version as XGBoost[1] throughout the paper. Since our implementation of XGBoost[1] failed to reproduce the performance metrics reported in [1], presumably due to undisclosed hyperparameters or training details, we implemented an additional XGBoost model with hyperparameter optimisation via Optuna. Apart from XGBoost, we further implemented random forest (RF) and support vector machine (SVM) algorithms with their default hyperparameters as baseline models.
The work proposing TabPFN benchmarks against gradient boosting methods, including XGBoost, across 18 numerical datasets, because gradient boosting methods continue to dominate tabular classification tasks [12]. We propose a privacy-aware framework to address the unavailability of raw videos due to privacy and ethics constraints. In such settings, tabular representations of pre-extracted visual features are often the only available modality. Prior research [1] on the HRI benchmark also established XGBoost as the best method, outperforming deep learning models. While advanced video models in this dataset could capture more expressive temporal cues, that would require access to full video sequences, which is unavailable to “protect privacy” [1]. Therefore, the above selection of baseline models is appropriate, given the tabular nature of the benchmark.
IV-C Results: Competitive performance on established evaluation protocol
Following the 5-fold stratified 10-repeats cross-validation evaluation protocol established by [1], Table II compares cross-validated average and standard deviation scores of ICL and fine-tuning of TFMPathy against the baseline models as well as the performance reported in the literature. Since TabPFN’s recommended input capacity is up to 500 features [13], we select 500 features. For a fair comparison, we also select 500 features for baseline models, including XGBoost, RF and SVM.
TFMPathy in ICL setup surpasses or achieves comparable performance to baseline models across all evaluation metrics. Assuming the scores reported for XGBoost[1] are average across 50 folds, the fine-tuned TabPFN still outperforms it in classification accuracy, precision and F1 scores. The discrepancy between the reported scores for XGBoost[1] and our re-implementation is likely due to undisclosed hyperparameters or training details.
TabICL outperforms TabPFN. Consistent with the findings of [17], our experiment confirms the superior performance of TabICL over TabPFN in the ICL setting. Note that we did not perform any hyperparameter tuning for the ICL experiments.
Fine-tuning tabular foundation model further improves performance. When fine-tuned, TabPFN shows a substantial performance gain and surpasses TabICL. This improvement, however, comes at a cost of additional computation. TabPFN has a total of 7.2 million trainable parameters, and in our experiment on a 64 GB GPU, the average fine‑tuning time per step ranged from 3.86 to 4.08 seconds for each of the 50 folds.
TFMPathy outperform other deep learning approaches. As shown in Table II, TFMPathy outperform LSTM and TCN deep learning architectures. Their lower performance may be attributed to the use of raw temporal sequences as input, rather than statistical attributes.
TABLE III: Generalisation on unseen subjects through leave-one-subject-out cross-validation. Highest scores are highlighted in bold. ICL refers to in-context learning, and FT refers to fine-tuning.
@@ IV-C Results: Competitive performance on established evaluation protocol
TFMPathyICL(TabICL) 0.541 0.519 0.547 0.475 0.509
TFMPathyICL(TabPFN) 0.459 0.472 0.451 0.377 0.411
TFMPathyFT(TabPFN; all layers trainable) 0.689 0.642 0.780 0.525 0.627
TFMPathyFT(TabPFN; frozen X and Y encoders) 0.730 0.669 0.818 0.590 0.686
IV-D Results: Improved generalisation on unseen subjects
Given the presence of multiple samples per subject (122 samples from 45 subjects), we adopt a train-test split that accounts for subject identity. Specifically, we use leave-one-subject-out cross-validation, where the splitting is based on participant ID available on the dataset. As this approach introduces no randomness, the results reported in Table III are deterministic and presented without any standard deviation.
Fine-tuning gives a significant boost to TabPFN cross-subject generalisation. Its classification accuracy improved from 0.459 to 0.730 after fine-tuning. Although the model had no access to the data from the test subject, it effectively learned patterns from other subjects, which suggests good cross-subject generalisation. For each of the 42 subjects, the average fine‑tuning time per step ranged from 4.70 to 4.98 seconds.
The primary architectural difference between TabPFN and TabICL lies in their attention mechanisms: TabICL uses a column-then-row attention scheme, which makes it faster than TabPFN. Since TabICL outperformed TabPFN in the ICL setup, it is reasonable to expect that fine-tuning TabICL could further improve performance, which we leave for future work.
@@ IV-D Results: Improved generalisation on unseen subjects
TABLE IV: Fine-tuning of TFMPathy(TabPFN) results (5-fold 10-repeats stratified subject-unaware cross-validation) with different experimental settings. Bold indicates the selected configuration.
@@
IV-E Ablation study
@@
Table IV reports a series of experiments on fine-tuning TabPFN (TFMPathyFT). Key hyperparameters explored include learning rate, batch size and optimiser choice. We tried the recent scheduler-free optimiser [37] as well as the standard AdamW optimiser, with and without a learning rate scheduler. Overall, TabPFN showed minimal performance variations across optimiser choices, with AdamW and one-cycle scheduler showing the best results. Like optimiser choice, batch size has a modest effect on performance. Increasing it from 8 to 32 improved classification accuracy from 0.701 to 0.708 and AUC from 0.689 to 0.696.
We also tested freezing three pre-trained layers of TabPFN: X encoder of 768 trainable parameters, Y encoder of 576 parameters, and Transformer encoder of 7.1 million trainable parameters. Freezing the X and Y encoders led to improved performance, likely because they are already well-generalised from TabPFN’s large-scale pre-training. Freezing preserves generalisable representations and enables faster fine-tuning, so we kept the X and Y encoders frozen and kept the remaining layers trainable.
@@
We further varied the softmax temperature and the number of input features. Using 500 features led to a slight performance improvement. A softmax temperature of 1.0 (i.e., no temperature scaling) yielded the best classification accuracy (Table IV). Based on these findings, the final model uses the AdamW optimiser with a one-cycle scheduler, a batch size of 32, frozen X and Y encoders, a softmax temperature of 1.0, and 500 selected input features.
@@
Figure 3: In-context learning performance of TFMPathy across different proportions of in-context data (i.e., training data used in the single forward pass of in-context learning). Results are based on 5-fold, 10-repeat stratified cross-validation without subject group separation.
@@
Figure 3 shows the ICL performance of the TFMPathy system using both TabICL and TabPFN across varying amounts of context data. In the ICL setup, the context data – i.e., training data – is fed into the model during the single forward pass. Both models generally exhibit improved performance as the amount of context data increases, which suggests the potential for further gains with additional data. However, TabPFN’s AUC peaked when using 80% of the available context data.
IV-F Limitations